The Bob of a simple pendulum oscillates with an amplitude of 20 cm. If the acceleration of the Bob is - 4y, where y is the displacement, calculate the



(I) magnitude of its maximum acceleration.



(II) period of the motion



(III) magnitude of the maximum velocity

1
Expert's answer
2022-04-06T14:03:27-0400

Given:

"y_m=0.2\:\rm m"

"a=-4y"


The acceleration of pendulum's bob

"a=-\omega^2y"

Hence, the angular frequency

"\omega=2\:\rm rad/s"

(I) magnitude of its maximum acceleration

"a_m=\omega^2 y_m=4*0.2=0.8\:\rm m/s^2"

(II) period of the motion

"T=\frac{2\pi}{\omega}=\frac{2\pi}{2}=\pi\:\rm s=3.14\: s"

(III) magnitude of the maximum velocity

"v_m=\omega y_m=2*0.2=0.4\:\rm m/s"